A Ferris wheel rotate 3 times each minute. the passengers sit in seats that are 25 feet from the center of the wheel.what is the angular velocity of the wheel in degrees per minute and radians per minute?what is the linear velocity of the passengers in the seats?
1
Expert's answer
2011-11-22T12:51:54-0500
L=400/4=100
L=2*pi*R
R=L/(2*pi)=100/6.28=15.92
w=2*pi*4=6.28*4=25.12